What is probability range for a distribution like this if it is equally likely that women or URM would be hired by either mechanism?

Agency-Wide Hires IC Search Hires
Male (Non-URM) 47 75
Female (Non-URM) 35 35
Female (URM) 8 2
Male (URM) 5 4
Total 95 116
Answer #1

Number of female or URM = Total - (male and non-URM) = (95+116) - (47+75) = 211 - 122 = 89

So, probability = 89/211 = 0.4218